# Evidence-Decision-Feedback: Theory-Driven Adaptive Scaffolding for LLM Agents

### v1

@@ Abstract
Multi-agent LLM architectures offer opportunities for pedagogical agents to help students construct domain knowledge and develop critical-thinking skills, yet many operate on a “one-size-fits-all” basis, limiting their ability to provide personalized support. To address this, we introduce Evidence-Decision-Feedback (EDF), a theoretical framework for adaptive scaffolding using LLMs. EDF integrates elements of intelligent tutoring systems and agentic behavior by organizing interactions around evidentiary inference, pedagogical decision-making, and adaptive feedback. We instantiate EDF through Copa, an agentic collaborative peer agent for STEM+C problem-solving. In an authentic high school classroom study, we show that EDF-aligned interactions align feedback with students’ demonstrated understanding and task mastery; promote gradual scaffold fading; and support interpretable, evidence-grounded explanations without fostering overreliance.
@@ 1 Introduction
Large language model (LLM) agents afford dialogic support that aligns with learning theories emphasizing knowledge construction through social interaction, including Social Cognitive Theory [1] (SCT), social constructivist perspectives, and the Zone of Proximal Development (ZPD) [30]. However, recent studies have examined the extent to which LLMs support (or hinder) learning, with emerging findings raising concerns: Kosmyna et al. [18] reported 83% of students receiving ChatGPT assistance during essay writing were unable to quote their own work, and Zhou et al. [32] found GPT-4 feedback on compiler errors improved performance and persistence during assisted tasks but not after feedback was withdrawn. These agents can encourage overreliance, thereby impeding sustained learning and critical thinking. As such, Borchers et al. [3] concluded that LLM tutoring in its current form is “unlikely to produce learning benefits rivaling known-to-be-effective ITS tutoring.”
@@ 2 Background
With agentic and multi-agent architectures, decision-making can be modularized, and interactions among tools and agents can be traced back to underlying decision-making processes. Li et al.’s EduMAS framework [19] employs Concept Clarifier and Experimentation Mentor agents that collaborate to support comprehensive learning experiences. Chu et al.’s GradeOpt system [4] integrates Grader, Reflector, and Refiner agents to improve alignment with human experts. Dai et al.’s Agent4EDU framework [12] uses agentic workflows to enable targeted instructional support. However, these systems provide limited guidance on grounding agent scaffolding in pedagogy.
Many existing agents also offer limited adaptive dialogic support. Google’s LearnLM-Tutor provides scaffolding informed by stakeholder co-design [17] but does not personalize its interactions. Thomas et al. [29] ground system scaffolding in feedback theory but without dialogic interaction. Other studies similarly employ LLMs to deliver single-shot feedback, lacking the interactive and adaptive qualities central to effective dialogic learning [23, 31, 20].
@@ 3 Theoretical Framework
Over four years of participatory design (PD) with students, teachers, and learning scientists, we synthesized surveys, interviews, focus groups, pilot studies, and collaborative design activities into a framework aligned with established theory [10, 5, 9, 8, 11, 13]. We present Evidence-Decision-Feedback (EDF), a principled approach to adaptive scaffolding for LLM-based pedagogical agents that addresses the limitations of existing frameworks discussed in Section 2.
@@ 4 Copa
Copa reasons over student log and chat data, selects actions, and generates feedback while under constraints. This aligns with Jurenka et al.’s [17] “see what the student sees” agent design principle and prior work suggesting bounded autonomy [24]—which balances agent initiative with pedagogical guardrails, e.g., human-in-the-loop methods [13]—represents an ideal design point to support learning while preserving trust and interpretability [11, 8]. Rather than relying on unconstrained long-horizon planning, Copa’s behavior is structured around the three EDF modules. Copa’s four-agent architecture is grounded in PD, which identifies appropriate agent behavior and the data to be considered, and in prior work demonstrating that modular, multi-agent systems improve transparency and task decomposition in educational settings [19, 4, 12].
@@
The StrategyAgent analyzes temporal action patterns to infer cognitive problem-solving behaviors, distinguishing productive strategies (e.g., TINKERING: systematic trial-and-error) from haphazard approaches (e.g., DEPTH-FIRST ENACTING: randomly building without periodic testing). The AssessmentAgent triangulates the current model state and strategy with past learner states (e.g., DEBUGGING or STRUGGLING) to generate an updated learner state. The KnowledgeAgent compares the model state to an expert reference to identify conceptual gaps within the ZPD and retrieves aligned domain knowledge (i.e., kinematics/computing information) via retrieval-augmented generation (RAG).
@@ 6 Discussion and Conclusions
Theory emphasizes that effective support should help students build understanding rather than just providing answers. However, consistent with prior work [7], our findings indicate that students do not always value this approach. Research also emphasizes the importance of learner agency for engagement [5], raising questions about the appropriate degree of autonomy when it conflicts with established pedagogical principles. Authors et al. [7] argue that student agency should be supported only to the point that it does not undermine critical thinking opportunities, a view supported by our findings. Overall, our results indicate that while EDF allows Copa to apply learning theories effectively, they reveal a tension between principled guidance and students’ preferences.
@@
Practically, effective deployment of agentic systems in situ requires attention beyond system design. As Thomas et al. [29] note, the value of agent feedback ultimately depends on students’ willingness to engage with it. Our findings reinforce the need for the AIED community to invest in fostering student buy-in and trust, as meaningful learning benefits are unlikely if students disengage from or resist agent support. Deeper engagement is, therefore, a prerequisite for realizing the pedagogical potential of LLM agents in practice.

### v2

@@ Abstract
Multi-agent LLM architectures offer opportunities for pedagogical agents to help students construct domain knowledge and develop critical-thinking skills, yet many operate on a “one-size-fits-all” basis, limiting their ability to provide personalized support. To address this, we introduce Evidence-Decision-Feedback (EDF), a theoretical framework for adaptive scaffolding using LLMs. EDF integrates elements of intelligent tutoring systems and agentic behavior by organizing interactions around evidentiary inference, pedagogical decision-making, and adaptive feedback. We instantiate EDF through Copa, an agentic collaborative peer agent for STEM+C problem-solving. In an authentic high school classroom study, we show that EDF-guided interactions align feedback with students’ demonstrated understanding and task mastery; promote gradual scaffold fading; and support interpretable, evidence-grounded explanations without fostering overreliance.
@@ 1 Introduction
Large language model (LLM) agents afford dialogic support that aligns with learning theories emphasizing knowledge construction through social interaction, including Social Cognitive Theory [1] (SCT), social constructivist perspectives, and the Zone of Proximal Development (ZPD) [30]. However, recent studies have examined the extent to which LLMs support (or hinder) learning, with emerging findings raising concerns: Kosmyna et al. (2025) [18] reported 83% of students receiving ChatGPT assistance during essay writing were unable to quote their own work, and Zhou et al. (2025) [32] found GPT-4 feedback on compiler errors improved performance and persistence during assisted tasks but not after feedback was withdrawn. These agents can encourage overreliance, thereby impeding sustained learning and critical thinking. As such, Borchers et al. (2025) [3] concluded that LLM tutoring in its current form is “unlikely to produce learning benefits rivaling known-to-be-effective ITS tutoring.”
@@ 2 Background
With agentic and multi-agent architectures, decision-making can be modularized, and interactions among tools and agents can be traced back to underlying decision-making processes. Li et al.’s (2024) EduMAS framework [19] employs Concept Clarifier and Experimentation Mentor agents that collaborate to support comprehensive learning experiences. Chu et al.’s (2025) GradeOpt system [4] integrates Grader, Reflector, and Refiner agents to improve alignment with human experts. Dai et al.’s (2025) Agent4EDU framework [12] uses agentic workflows to enable targeted instructional support. However, these systems provide limited guidance on grounding agent scaffolding in pedagogy.
Many existing agents also offer limited adaptive dialogic support. Google’s LearnLM-Tutor provides scaffolding informed by stakeholder co-design [17] but does not personalize its interactions. Thomas et al. (2025) [29] ground system scaffolding in feedback theory but without dialogic interaction. Other studies similarly employ LLMs to deliver single-shot feedback, lacking the interactive and adaptive qualities central to effective dialogic learning [23, 31, 20].
@@ 3 Theoretical Framework
Over four years of participatory design (PD) with students, teachers, and learning scientists, we synthesized surveys, interviews, focus groups, pilot studies, and collaborative design activities into a framework aligned with established theory [10, 5, 7, 9, 8, 11, 13]. We present Evidence-Decision-Feedback (EDF), a principled approach to adaptive scaffolding for LLM-based pedagogical agents that addresses the limitations of existing frameworks discussed in Section 2.
@@ 4 Copa
Copa reasons over student log and chat data, selects actions, and generates feedback while under constraints. This aligns with Jurenka et al.’s (2024) [17] “see what the student sees” agent design principle and prior work suggesting bounded autonomy [24]—which balances agent initiative with pedagogical guardrails, e.g., human-in-the-loop methods [13]—represents an ideal design point to support learning while preserving trust and interpretability [11, 8]. Rather than relying on unconstrained long-horizon planning, Copa’s behavior is structured around the three EDF modules. Copa’s four-agent architecture is grounded in PD, which identifies appropriate agent behavior and the data to be considered, and in prior work demonstrating that modular, multi-agent systems improve transparency and task decomposition in educational settings [19, 4, 12].
@@
The StrategyAgent analyzes temporal action patterns to infer cognitive problem-solving behaviors, distinguishing productive strategies (e.g., TINKERING: systematic trial-and-error) from haphazard approaches (e.g., DEPTH-FIRST ENACTING: randomly building without periodic testing). The AssessmentAgent triangulates the current model state and strategy with past learner states (e.g., EXPLORING, DEBUGGING) to generate an updated learner state. The KnowledgeAgent compares the model state to an expert reference to identify conceptual gaps within the ZPD and retrieves aligned domain knowledge (i.e., kinematics/computing information) via retrieval-augmented generation (RAG).
@@ 6 Discussion and Conclusions
Theory emphasizes that effective support should help students build understanding rather than just providing answers. However, consistent with prior work [7], our findings indicate that students do not always value this approach. Research also emphasizes the importance of learner agency for engagement [5], raising questions about the appropriate degree of autonomy when it conflicts with established pedagogical principles. Authors et al. (2025) [7] argue that student agency should be supported only to the point that it does not undermine critical thinking opportunities, a view supported by our findings. Overall, our results indicate that while EDF allows Copa to apply learning theories effectively, they reveal a tension between principled guidance and students’ preferences.
@@
Practically, effective deployment of agentic systems in situ requires attention beyond system design. As Thomas et al. (2025) [29] note, the value of agent feedback ultimately depends on students’ willingness to engage with it. Our findings reinforce the need for the AIED community to invest in fostering student buy-in and trust, as meaningful learning benefits are unlikely if students disengage from or resist agent support. Deeper engagement is, therefore, a prerequisite for realizing the pedagogical potential of LLM agents in practice.